OpenStudy (anonymous):

What is the vertex of the graph of y = −4(x + 2)2 + 5?

OpenStudy (mathmale):

this equation is already in the form of the equation of a parabola that shows the coordinates of the vertex: y = a(x-h)^2 + k y =-4 (x-[-2])^2 + k compare the two equations. What do you think is the x-coordinate of the vertex? The y-coordinate?